OpenStudy (anonymous):

Multiply and simplify by factoring sqrt5b^3 sqrt55c^4

OpenStudy (anonymous):

\[\sqrt{5b ^{3}} \sqrt{55c ^{4}}?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[5bc ^{2}\sqrt{11b}\]